# Distinguishing Neutron Star vs. Low-Mass Black Hole Binaries with Late Inspiral & Postmerger Gravitational Waves — Sensitivity to Transmuted Black Holes and Non-Annihilating Dark Matter

### v2

@@ Abstract
The astrophysical origin of observed low-mass compact binary coalescences in the $1-2.5\,M_{\odot}$ range remains ambiguous. Both binary neutron star (BNS) and binary low-mass black hole (LMBH) mergers produce nearly identical inspiral waveforms, and electromagnetic follow-up is not always possible. Distinguishing between these scenarios therefore presents a key challenge. We demonstrate that waveform differences in the late-inspiral to postmerger epochs create significant mismatches that will be detectable by planned detectors, viz., NEMO, Cosmic Explorer, and Einstein Telescope, while the currently operational LIGO A+ will be effective only for nearby sources. These differences are enhanced for stiffer equations of state. We show how the redshift-dependent compact binary merger rate inferred from gravitational wave observations can be parsed into BNS and LMBH components, accounting for misclassification probability. We forecast model-independent 90% exclusion sensitivities for the LMBH fraction. Interpreting these LMBHs as dark matter capture-induced transmuted black holes, we convert exclusion sensitivities into projected exclusion bounds on heavy non-annihilating dark matter. Our results illustrate how gravitational wave measurements can disentangle compact object populations and provide new insights into particle dark matter interactions.
@@ 1 Introduction
Gravitational wave (GW) [1] observations of compact binary mergers provide key insights into strong-field gravity and relativistic astrophysics. GW170817 [2], the first observed binary neutron star (BNS) merger, enabled multimessenger astronomy by linking gravitational waves to electromagnetic counterparts. GW150914 [3], the first detected binary black hole (BBH) merger, confirmed the existence of stellar-mass black hole (BH) binaries and provided a new way to study general relativity.
@@
During the O3a+b observing runs, the LIGO-Virgo-KAGRA (LVK) Collaboration [4] has reported a number of compact binary coalescence (CBC) events with component masses that resist clear classification as either neutron stars (NSs) or low-mass black holes (LMBHs). Examples include GW190425 [5], GW190814 [6], GW230529 [7], and SSM200308 [8], where at least one of the components is either in the lower mass-gap region, viz. $2.5-5\,M_{\odot}$, or even lighter [9, 10, 11]. Such LMBHs are not expected as a product of standard stellar evolution, and thus the current practice is to provisionally classify CBCs as NSs or BHs using the inferred mass. Nevertheless, such a classification lacks robustness, and a more comprehensive approach is warranted.
@@
In this paper, we therefore ask the question — Can we distinguish BNSs and LMBHs? This is motivated from two complementary viewpoints. Firstly, from an astrophysical/observational standpoint, it is desirable to classify a low-mass CBC as a BNS or an LMBH (we assume that these are the only two possibilities). This is important in order to obtain unbiased inferences on the NS population and merger rates. Secondly, from a fundamental physics perspective, the potential existence of these exotic LMBHs offers novel avenues to discover new physics [12, 13, 14, 15].
@@
A specific possibility involves dark matter (DM) [16, 17], which constitutes  27% of the Universe’s energy content. Despite its ubiquity, its particle identity remains elusive — with its mass uncertain within a huge range ($10^{-22}-10^{66}\,\rm eV$) [18] and no confirmed non-gravitational interactions. Non-annihilating DM particles, e.g., with masses in the range $10^{2}-10^{7}$ GeV for bosons, are capable of producing LMBH mergers that mimic BNS signals. Galactic DM with non-gravitational nucleon interactions can be captured by NSs over their lifetimes [19, 20, 21, 22, 23, 24, 14, 25, 26, 27, 16, 28, 29]. Once captured, heavy DM can accumulate in the NS core, forming a dense dark core that may become gravitationally unstable and collapse into a microscopic BH. This BH can eventually consume the entire NS, ejecting $\sim 10^{-3}-10^{-4}\,M_{\odot}$[30], and forming an LMBH of comparable mass [31, 22, 23, 32, 14, 33, 34]. Thus, a BNS system can evolve into an LMBH–LMBH system via DM-induced collapse. Such mergers should, in principle, be detectable by current GW instruments. Their apparent absence places constraints on DM mass and nucleon cross-sections [28], though these are tentative since the LVK collaboration classifies all compact objects below $2.5\,M_{\odot}$ as NSs [35]. Developing robust methods to identify such objects is therefore essential. Apart from this scenario, GW observations have been used to study DM-admixed binary neutron star mergers [36, 37], infer DM-induced modifications to NS oscillation modes [38, 39], and probe DM density spikes around massive black holes using extreme mass ratio inspirals (EMRIs) [40, 41, 42, 43]. In addition, GW observations have also shed light on possible nature of DM by probing dark photons as DM [44], ultralight scalars [45], non-annihilating DM [28, 46, 47], compact DM [48, 49, 50, 51], and self-interacting DM [52, 53], among others. It has been well established that GW observations can serve as a nobel probe of DM [54], complementing traditional direct, indirect and collider-based searches.
@@
Previous work on distinguishing BNSs and LMBH mergers has considered the potential of multi-messenger detection. As in the case of GW170817 [2], non-gravitational signals would provide clear evidence against LMBHs. However, it is estimated that most of the CBC detections at current and future GW detectors will not be adequately followed up by electromagnetic and neutrino searches [55, 56, 57]. This has to do with the limited sky coverage and/or angular resolution of the relevant telescopes, as well as imperfect localization of the GW sources.
@@
In the absence of an electromagnetic or neutrino counterparts, it is challenging to distinguish a BNS merger from an LMBH merger, as the current GW observations primarily capture the inspiral phase, where the waveforms of these systems are nearly identical. This degeneracy becomes especially relevant for merger events with component masses in the $2-5\,M_{\odot}$ range. One promising approach involves measuring tidal effects during the inspiral. Neutron stars, being extended objects with internal structure, exhibit non-zero tidal deformability; whereas black holes, being point-like singularities without internal structure, have zero tidal polarizability. Accurate inference of the tidal parameters in the waveform thus offers a potential diagnostic to distinguish BNS from LMBH mergers [47]. Recent studies have explored this avenue [58, 59, 60, 61, 62, 63, 64, 65, 66, 67], although such analyses require higher order post-Newtonian corrections (up to 5PN), adding significant computational complexity. Further, these tidal effects during the inspiral are subtle and typically do not offer any robust qualitative features, independent of the neutron star equation of state (EoS) that remains uncertain due to possible exotic matter at supranuclear densities [68, 69]. Thus, reliable extraction of tidal signatures demands favorable observational conditions: short luminosity distances, long inspiral durations with high signal-to-noise ratios (SNRs), and enhanced detector sensitivity at high frequencies (up to 343-4 kHz), since the merger frequencies for $_{\odot}-1M1\,M_{\odot}$ systems are approximately 2 kHz [70].
@@
To address the question posed at the beginning, in this paper we systematically investigate the similarity between GW waveforms from equal-mass LMBH–LMBH and BNS–BNS mergers, by comparing LMBH-LMBH mergers waveforms from numerical relativity simulations of the late-inspiral and postmerger epochs. We compute the Fitting Factors and Bayes Factors between these waveforms and identify the conditions under which they can be distinguished. To evaluate observational prospects, we analyze the sensitivity of both current and future GW detectors, viz., LIGO A+ [71], NEMO [72], CE [73], and ET [74]. The main outcome of these studies is an estimate of the efficiency of correctly classifying a CBC as a BNS or an LMBH. We use this efficiency factor to estimate the most probable BNS and LMBH fraction in the CBC event rate, as well as to obtain exclusion sensitivity for the fraction of CBCs that could be LMBHs. Finally, we explore a physically motivated model for formation of LMBHs, via transmutation of NSs into similar mass BHs due to capture of heavy non-annihilating DM particles, and compute exclusion sensitivity to the DM mass $m_{\chi}$ and its interaction cross section $\sigma_{\chi n}$ with nucleons.
The paper is organized as follows: In Sec. 2, we introduce the benchmark waveforms of LMBH-LMBH and NS-NS mergers using the PyCBC and CoRe simulation frameworks. Sec. 3 presents the computation of the Fitting Factors and Bayes Factors, its dependence on EoSs, detector characteristics, and the source luminosity distance. In Sec. 4, we present the results on BNS and LMBH merger rates and on DM parameters, and conclude in Sec. 5 with a brief summary and outlook.
2 Inputs: LMBH & BNS Waveforms and Benchmarks
2.1 LMBH Waveforms from PyCBC
[…]
Low-mass black holes have become a new avenue to look for the BSM physics through gravitational wave observations. In this section we will introduce the waveform of a LMBH merger, irrespective of its origin and evolution. We consider two non-spinning LMBHs with equal masses, $m_{1}=m_{2}=1.35M_{\odot}$, for which the chirp mass is $M_{c}=\mu^{3/5}\,M^{2/5}=1.17M_{\odot}$, where $\mu\equiv{m_{1}m_{2}}/(m_{1}+m_{2})$ is the reduced mass of the system and $M\equiv\left(m_{1}+m_{2}\right)$ is the total mass of the system. The ‘plus’ (+) and ‘cross’ ($\times$) polarized strain amplitudes for the $(l=2,m=2)$-mode can be found as [70],
$\tilde{h}_{\times}(f)=\left(\frac{5}{24}\right)^{1/2}\frac{1}{\pi^{2/3}}\frac{1}{D_{L}}\left(GM_{c}\right)^{5/6}f^{-7/6}e^{i\Psi_{\times}(f)}\cos{\iota},$ (2.1)
where $G$ is the gravitational constant, $D_{L}$ is the luminosity distance, $M_{c}$ is the detector-frame chirp mass given by $(1+z)M_{c}|_{\rm source\,rest\,frame}$. The redshift $z$ can be assumed to be 0 for current observations so far. The inclination of the perpendicular to the orbital plane with respect to the line of sight is given as $\iota$}. If $\iota=\pi/2$, the dominant strain amplitude is reduced by a factor of 1/2 and $h_{\times}$ vanishes, resulting in a linearly polarized wave; this orientation is termed as ‘edge-on’. For $\iota=0\,\emph{\&}\,\pi$, the system’s orbital angular momentum is aligned with the line of sight and this is termed as a ‘face-on’ orientation, resulting in a circularly polarized wave. For a face-on symmetric $\left(m_{1}=m_{2}\right)$ LMBH binary at $D_{L}=$ 1 Mpc (with vanishing spins and $m_{1}=m_{2}=1.35\,M_{\odot}$), the strain amplitude will be
$\tilde{h}(f)\approx 4.2\times 10^{-22}\,{\rm sec}\left(\frac{\rm Mpc}{D_{L}}\right)\left(\frac{M_{c}}{1.17M_{\odot}}\right)^{5/6}\left(\frac{f}{100\,\rm Hz}\right)^{-7/6}\,.$ (2.2)
This waveform is valid up to a maximum frequency, termed as the frequency of innermost stable circular orbit (ISCO), given by
$f_{\rm ISCO}=\frac{1}{6\sqrt{6}\left(2\pi\right)}\frac{c^{3}}{GM}\approx 2.2\,{\rm kHz}\,\left(\frac{M_{\odot}}{M}\right)\,.$ (2.3)
To generate the LMBH waveforms we use the publicly available software PyCBC [79]. We employ its time domain and frequency domain module named IMRPhenomD [75, 76], which is a phenomenological model of GW signal from the inspiral-merger-ringdown of non-precessing (aligned spin) black hole binaries. These waveforms are generated using numerical relativity (NR) simulations and cover the entire coalescence, including the strong gravity regimes of merger and ringdown unlike the waveforms generated using perturbative expansion in $v/c$, where $v$ is the orbital velocity and $c$ is the speed of light. We show symmetric LMBH binary frequency domain waveforms in Fig. 1 with the total masses $2M_{\odot}$, $2.7M_{\odot}$, and $4M_{\odot}$ respectively. These assume non-spinning BHs and starts with a frequency of 40 Hz with frequency steps $\left(\Delta f\right)$ of 0.125 Hz, with $D_{L}=1$ Mpc.
2.2 BNS Waveforms from CoRe Database & Choice of EoSs
Binary neutron star mergers, observed through gravitational waves and electromagnetic signals, provide key insights into multi-messenger and particle astrophysics, as confirmed by GW170817 [2]. Whereas, events like GW190425 [5] and GW190814 [6] open a new window for exotic physics and constrain the NS EoSs.
[…]
CoRe provides a diverse set of GW templates to constrain the neutron star equation of state, and to analyze merger dynamics and remnant properties. The first release of CoRe [87] comprised 367 waveforms derived from numerical simulations adhering to general relativity principles and, 164 distinct setups, varying parameters such as total binary mass, mass ratio, initial separation, eccentricity, and stellar spins. The second release [88] included 254 distinct BNS configurations, resulting in a total of 590 numerical-relativity simulations performed at various grid resolutions. These simulations encompass a wide range of possibilities with 18 different EoSs and, with a wide range of parameters. These simulations are executed with two independent mesh-based methods: BAM [89, 90] and THC [91]. BAM is better adapted to treat the NS surface, multiple orbits in the inspiral, followed by the merger and post-merger. Whereas THC is preferentially used to address the microphysics of the NS, as it contains different neutrino transport schemes, different mixing and dissipation mechanisms during the merger. For this work, we chose 6 EoSs from BAM simulations and 2 from THC simulations. The absolute value of the strain amplitude is more important for our study, than the phase evolution associated with the detailed microphysics.
$\chi_{\rm eff}=\frac{m_{1}\chi_{1}+m_{2}\chi_{2}}{M}\,$, (2.4)
$\tilde{\Lambda}=\frac{16}{13}\left[\frac{\left(m_{1}+12m_{2}\right)m_{1}^{4}\Lambda_{1}}{M^{5}}+\frac{\left(m_{2}+12m_{1}\right)m_{2}^{4}\Lambda_{2}}{M^{5}}\right]\,$, (2.5)
$h_{+}-ih_{\times}=D_{L}^{-1}\sum_{l=2}^{\infty}\sum_{m=-l}^{l}h_{lm}(t)_{\,-2}Y_{lm}(\iota,\phi)\,$, (2.6)
where $_{-2}Y_{lm}$ are the s=2s=-2 spin-weighted spherical harmonics [70]. The inclination and orbital phase, $\iota$ and $\phi$, respectively, give the inclination of the perpendicular to the orbital plane with respect to the line of sight, and the orbital phase of the binary at a particular time.
$P\left(\rho,\epsilon\right)=P_{\rm pwp}\left(\rho\right)+\left(\gamma_{\rm th}-1\right)\rho\left(\epsilon-\epsilon_{\rm pwp}\right)\,$, (2.7)
where $P_{\rm pwp}(\rho)$ is a given piecewise polytropic EoS model [94]. The specific parameters employed for the piecewise polytropic EoS are available on the CoRe website [80] and they mimic the zero temperature EoS models described in [94]. Fig. 1 of Ref. [88] shows the $m_{\rm NS}-R_{\rm NS}$ diagram and the $\Lambda-m_{\rm NS}$ diagram of these EoS models, which gives an overview of the current parameter space of EoS models. For a given mass, $M=1.4M_{\odot}$, the EoS 2H furnishes the largest radius $R^{\rm TOV}$ $\approx$ 15.21 km, and EoS 2B the smallest $\approx$ 9.75 km [88, 95]. Evidently 2H is the stiffest (high pressure for a given density) and 2B is the softest (lower pressure for a given density) EoS as used in this database. The larger (smaller) values of tidal polarizability parameter $\Lambda$ correspond to stiffer (softer) EoSs, resulting in long-lived remnant (prompt collapse) in the post-merger. In the following section, we outline the benchmark parameters adopted for this analysis and specify the set of EoSs selected for our study.
Figure 2: Frequency domain waveforms of $1.35M_{\odot}-1.35M_{\odot}$ BNS mergers, with non-spinning components, for the EoSs described in Table 2. The key feature is the postmerger second peak which may furnish a way to distinguish BNS from LMBHs.
The extraction of BNS waveforms from this database is discussed in detail (with the file structure and nomenclature) in the Appendix A.1. This data is in time-domain, and a Fourier transform gives the waveform in frequency domain. A sample of BNS waveforms from CoRe is shown in Fig. 2. They exhibit qualitative differences from that of a LMBH binary, primarily due to the matter effects. Depending on the EoS, the BNS merger remnant may result in a hypermassive NS that subsequently collapses to a BH, or undergo a prompt collapse. This behavior typically produces a characteristic second peak in the postmerger spectrum. Note the second peak seen for all the EoSs, at $f\lesssim 4$ kHz, except for 2B which is in tension with NS mass-radius data [88].
2.3 Benchmark Parameters
We now summarize the fiducial parameters used throughout our analysis. These include the intrinsic source properties, the NS EoSs adopted for BNS mergers, and the assumed source $D_{L}$ for different detector configurations.
Table 1 lists the intrinsic parameters used for both LMBH and BNS systems. In both cases, we consider equal-mass binaries with component masses $m_{1}=m_{2}=1.35\,M_{\odot}$ and vanishing dimensionless spin parameters $\chi_{1}=\chi_{2}=0$. For this study we have chosen eight EoSs, and generated the waveform in frequency domain. We use 6 BAM waveforms (2H, MS1b, H4, ALF2 [96], SLy, 2B) with zero-temperature EoSs and 2 finite-temperature EoSs ($\rm BHB\Lambda\phi$[97], LS220 [98]) for our study. We list them in Table 2 from stiffer to softer models [99, 88], along with their tidal coupling constant and the reduced tidal parameter (see Eq. 2.5), with their CoRe IDs alongside. Finally, Table 3 summarizes the source luminosity distances $D_{L}$, for each GW detector considered; LIGO A+ [71], NEMO [72], CE [73], and ET [74]. These distances are chosen based on the detector sensitivities (see Fig. 4), such that the norm remains significant for detection.
System Component Masses Dimensionless Spins
LMBH–LMBH $m_{1}=m_{2}=1.35\,M_{\odot}$ $\chi_{1}=\chi_{2}=0$
BNS–BNS $m_{1}=m_{2}=1.35\,M_{\odot}$ $\chi_{1}=\chi_{2}=0$
Table 1: Benchmark intrinsic parameters used in this study for LMBH and BNS systems. All sources are assumed to be equal-mass, non-spinning binaries.
EoS CoRe ID $\kappa_{2}^{T}$ $\tilde{\Lambda}$ $f_{\rm merge}$ in Hz
2H BAM:0002 436 2325 1230
MS1b BAM:0065 287 1531 1407
H4 BAM:0035 208 1111 1554
[…]
ALF2 BAM:0003 137 733 1743
LS220 THC:0019 128 684 1774
SLy BAM:0098 73 390 2001
2B BAM:0001 24 127 2298
Table 2: The EoSs, considered for this study, with their CoRe IDs. The reduced tidal parameters $\kappa_{2}^{T}$ & $\tilde{\Lambda}$ are also provided (Eq. 2.5). The last column gives the merger frequencies [99, 100, 101] for different EoSs, that separate the inspiral & postmerger phases.
Detector Source Luminosity Distance [Mpc]

### v3

@@ Abstract
Gravitational wave signals from binary neutron star (BNS) mergers and binary low-mass black hole (BLMBH) mergers are highly similar in the early inspiral phase. Consequently, the astrophysical origin of recently detected low-mass compact binary coalescences has remained ambiguous, particularly in the absence of electromagnetic counterparts. In this work, we demonstrate that proposed detectors with increased high-frequency sensitivity – including NEMO, Cosmic Explorer, and the Einstein Telescope – will reliably distinguish these two source classes in the late inspiral and postmerger regimes. We further show how these detections can be used to disentangle the individual contributions of BNS and BLMBH systems to the compact binary merger rate, while accounting for misclassification probabilities. Finally, we show this can lead to constraints on the interaction of heavy, non-annihilating dark matter with nucleons. This is achieved by noting that capture of such dark matter particles into neutron stars would lead to transmuted black holes (TBHs), formed via neutron star collapse, which would contribute to the BLMBH rate.
@@ 1 Introduction
Gravitational wave (GW) [1] observations of compact binary mergers provide key insights into strong-field gravity and relativistic astrophysics. While GW150914 [2], the first detected coalescence of two black holes (BHs), has confirmed the existence of stellar-mass binary black hole (BBH) mergers and provided a new way to study general relativity, GW170817 [3], the first observed merger of two neutron stars (NSs), has opened a new window to multimessenger astronomy by linking gravitational waves to electromagnetic counterparts. During the O3a+b observing runs, the LIGO-Virgo-KAGRA (LVK) Collaboration [4] has reported a number of compact binary coalescence (CBC) events with component masses that resist clear classification as either NSs or low-mass black holes (LMBHs). Examples include GW190425 [5], GW190814 [6], GW230529 [7], and SSM200308 [8], where at least one of the components is either in the lower mass-gap region, viz. $2.5-5\,M_{\odot}$, or even lighter [9, 10, 11]. Such LMBHs are not expected as a product of standard stellar evolution, and thus the current practice is to classify CBC components as NSs or BHs using the inferred mass (i.e. NS if $\lesssim 2.5\,M_{\odot}$, and BH if not). However, such a classification is provisional and a more rigorous data-driven approach is warranted.
@@
In this paper, we therefore ask the question — How well can we distinguish NS-NS and LMBH-LMBH mergers? This is motivated from two complementary viewpoints. Firstly, from an astrophysical/observational standpoint, it is desirable to classify a low-mass CBC as a binary neutron star (BNS) or a binary low-mass black hole (BLMBH) merger, assuming these are the only two possibilities. This is important in order to obtain unbiased inferences on the NS population and merger rates. Secondly, from a fundamental physics perspective, the potential existence of these exotic LMBHs offers a novel gravitational avenue to discover new physics [12, 13, 14, 15, 16].
@@
A specific new physics possibility involves dark matter (DM) [17, 18], which constitutes  27% of the Universe’s energy content. Despite its ubiquity, its particle identity remains elusive – with its mass uncertain within a huge range ($10^{-22}-10^{66}\,\rm eV$) [19] and no confirmed non-gravitational interactions. Heavy non-annihilating DM particles are capable of producing LMBHs via DM capture-induced collapse of a NS – i.e., a ‘transmuted black hole’ (TBH); we discuss this in more detail in Sec. 3.2. Interestingly, mergers of these TBHs can mimic BNS merger signals [16, 14], and their apparent absence places constraints on DM mass and nucleon cross sections [16, 20]. More broadly, GW observations can serve as a novel probe of DM [21], complementing traditional direct, indirect and collider-based searches in a variety of ways. [^1] Other related works connecting DM with GW observations include the exploration of DM-admixed BNS mergers [22, 23, 24], inferring DM-induced modifications to NS oscillation modes [25, 26], and probing DM density spikes around super-massive black holes [27] using extreme mass ratio inspirals (EMRIs) [28, 29, 30, 31]. In addition, GW observations have also shed light on the possible nature of DM by probing dark photons as DM [32], ultralight scalars [33], non-annihilating DM [16, 34, 35], compact DM [36, 37, 38, 39], and self-interacting DM [40, 41, 42], among others.
@@
Coming back to the issue of distinguishing BNS and BLMBH mergers, a key distinguishing feature between them is the potential observation of an electromagnetic counterpart through multi-messenger detection. As in the case of GW170817 [3], non-gravitational signals would provide clear evidence against BLMBHs. However, it is estimated that most of the CBC detections at current and future GW detectors will not be adequately followed up by electromagnetic and neutrino searches [43, 44, 45]. This has to do with the limited sky coverage and/or angular resolution of the relevant telescopes, as well as imperfect localization of the GW sources.
@@
In the absence of electromagnetic or neutrino counterparts, it is challenging to confidently distinguish a BNS merger from a BLMBH merger. Current GW detectors primarily record the inspiral phase signal, where the waveforms of these systems are nearly identical. A promising approach involves measuring tidal effects during the inspiral — Neutron stars, being extended objects with internal structure, exhibit non-zero tidal deformability; whereas black holes, being point-like singularities without internal structure, have zero tidal polarizability. Accurate inference of the tidal parameters in the waveform during the inspiral phase thus offers a potential diagnostic to distinguish BNS from BLMBH mergers [35, 46]. Several recent studies have explored this avenue [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. However, it is important to note that tidal effects can be degenerate with other effects such as spin or orbital eccentricity, unless high-order post-Newtonian effects are reliably modeled and detected [57]. A reliable extraction of tidal signatures demands favorable observational conditions: nearby sources, long inspiral durations with high signal-to-noise ratios (SNRs), and enhanced detector sensitivity at the higher end of currently detectable frequencies (note that the merger frequencies for $_{\odot}-1M1\,M_{\odot}$ systems are approximately 2 kHz) [58]. In contrast, matter effects in neutron stars become more pronounced and qualitatively distinct in the postmerger phase [59, 60, 46].
To address the question posed at the beginning, we therefore investigate the similarity between similar-mass NS-NS and LMBH–LMBH mergers, by comparing their waveforms in the late inspiral and postmerger epochs. We assume that the true events are in fact BNS mergers, and ask how well such events can be fit by some BLMBH waveform (see Sec. 2.2). This is then statistically quantified using fitting factors and Bayes factors, and we predict the probability to misclassify a true BNS merger as a BLMBH merger (see Sec. 2.3).
@@
The main outcome of these studies is an estimate of the probability to correctly classify a CBC as a BNS or a BLMBH. The results depend on the equation of state (EoS) of the neutron star matter and on the detector sensitivity. We then use this to estimate the most probable BNS and BLMBH contributions to the observed CBC rate, as well as to obtain exclusion sensitivity for the fraction of CBCs that could be BLMBHs (see Sec. 3.1). Finally, we explore a physically motivated model for formation of LMBHs, i.e. via transmutation of NSs into similar mass BHs due to capture of heavy non-annihilating DM particles, and compute exclusion sensitivity to the DM mass $m_{\chi}$ and its interaction cross section $\sigma_{\chi n}$ with nucleons (see Sec. 3.2).
Figure 1: Dimensionless strain amplitudes $f\tilde{h}(f)$ for a benchmark BNS merger and BLMBH merger (grey lines), and the amplitude spectral density (ASD) of the detector noise for the set of current and proposed detectors, LIGO A+ [61], NEMO [62], CE [63], and ET [64], considered in this study. Here the source is taken at a luminosity distance $D_{L}=1$ Mpc. The different ASDs are shown as $\sqrt{fS_{n}(f)}$ to keep it dimensionless. $S_{n}(f)$ denotes the one-sided power spectral density of the detector noise.
2 Distinguishability of BNS & BLMBH Mergers
In Fig. 1 we show the dimensionless strains for example symmetric BNS and BLMBH mergers, in comparison to the noise sensitivity curve of each of the detectors: LIGO A+ [61], NEMO [62], CE [63], and ET [64]. See Appendix A.3 for details of each of these detectors. The key observation is that, while at lower frequencies ($\lesssim 2\,$kHz) the waveforms are qualitatively similar (though quantitatively different), at higher frequencies the BNS merger produces a distinctive second peak and the signals have qualitative differences.
To quantify the distinguishability of the BNS and BLMBH waveforms, we will treat the BNS waveform for a given EoS as the putative detected waveform and fit it with BLMBH merger waveforms. The overlap between two waveforms is termed as match $\mathcal{M}$. The maximal match, obtained by scanning over the intrinsic parameters of BLMBH templates, is termed as the fitting factor $\rm FF$, which can then be used to estimate a Bayes factor BF. In the following, we first list the benchmark parameters and EoSs for our study, our procedure for calculating the FFs and BFs, and then compute the detector-specific distinguishing ability for the chosen set of EoSs.
2.1 Inputs & Benchmark Parameters
As our benchmark BNS waveforms, we use a set of 8 numerical relativity (NR) waveforms from the CoRe database [65]. Of these, 6 are BAM waveforms with zero-temperature EoSs (2H, MS1b, H4, ALF2, SLy, 2B) and 2 are finite-temperature THC EoSs ($\rm BHB\Lambda\phi$ and LS220). See Appendix A.2 for details of the CoRe database, EoSs, and practicalities of extracting the waveforms. All of these correspond to a merger of two non-spinning NSs, each with mass $1.35\,M_{\odot}$, in a non-eccentric orbit.
As our template bank, we use BLMBH merger waveforms generated using the IMRPhenomD [66, 67] model in PyCBC [68]. In this study, we restrict ourselves to equal-mass binaries, with each component mass in the range $1-2.5\,M_{\odot}$. We also assume that they are non-spinning and the orbits are not eccentric. See Appendix A.1 for further details.
(a) Intrinsic Parameters
BNS BLMBH
$m_{1}=m_{2}$ $1.35\,M_{\odot}$ $1-2.5\,M_{\odot}$
$\chi_{1}=\chi_{2}$ 0 0
Microphysics See Table 2
(b) Distances
Detector $D_{L}$ [Mpc]